A 4% salt solution is mixed with a 16% salt solution. How many milliliters of each solution are needed to obtain 600 millilters of a 10% solution?
Answer by nerdybill(7384)

A 4% salt solution is mixed with a 16% salt solution. How many milliliters of each solution are needed to obtain 600 millilters of a 10% solution?
.
Let x = amount of 4% salt solution
then
600-x = amount of 16% salt solution
.
.04(x) + .16(600-x) = .10(600)
.04x + 96 - .16x = 60
96 - .12x = 60
96 - 60 = .12x
36 = .12x
300 milliliters of 4% salt solution = x
.
600-x = 600-300 = 300 milliliters of 16% salt solution